# OViP: Online Vision-Language Preference Learning

### v1

@@ Abstract
Large vision-language models (LVLMs) remain vulnerable to hallucination, often generating content misaligned with visual inputs. While recent approaches advance multi-modal Direct Preference Optimization (DPO) to mitigate hallucination, they typically rely on predefined or randomly edited negative samples that fail to reflect actual model errors, limiting training efficacy. In this work, we propose an Online Vision-language Preference Learning (OViP) framework that dynamically constructs contrastive training data based on the model’s own hallucinated outputs. By identifying semantic differences between sampled response pairs and synthesizing negative images using a diffusion model, OViP generates more relevant supervision signals in real time. This failure-driven training enables adaptive alignment of both textual and visual preferences. Moreover, we refine existing evaluation protocols to better capture the trade-off between hallucination suppression and expressiveness. Experiments on hallucination and general benchmarks demonstrate that OViP effectively reduces hallucinations while preserving core multi-modal capabilities.
@@ 1 Introduction
Large vision-language models (LVLMs) [1, 2, 3, 4, 5] have demonstrated remarkable performance across a wide range of multi-modal tasks [6, 7, 8, 9] by integrating pre-trained visual encoders with large language models (LLMs) to process and generate language grounded in visual inputs. However, LVLMs continue to struggle with persistent hallucination issues [10, 11], often exhibiting incorrect references to visual content [12, 13, 11]. These errors manifest as misattributing object properties, describing nonexistent entities, or fabricating spatial relationships that do not align with the image. Such inconsistencies undermine the model’s faithfulness to the input and hinder further reasoning capabilities, significantly limiting the reliability of LVLMs in real-world applications.
@@
Recent success of Direct Preference Optimization (DPO) [14] in LLMs alignment has inspired the exploration of multi-modal DPO to mitigate hallucination in LVLMs [15, 16, 17, 18]. However, early efforts directly extend the original DPO designs from LLMs to LVLMs by constructing preference pairs solely on textual responses given the same image input, primarily focusing on response-side preference optimization and showing limited effectiveness. Recent advancements incorporate additional preference pairs conditioned on varying image inputs while keeping the same response, optimizing both visual and textual preference optimization [19, 20, 21]. This paradigm provides a complementary training signal that encourages the model to attend more closely to visual content.
@@
However, prior work mainly relies on existing paired datasets [20] or expert-defined pattens to construct negative image inputs, using techniques such as random cropping [19], noise disruption [22], object removal [23], or human/LLMs generated element-replaced response for image editing [17]. These strategies are typically not explicitly tied to model failures, resulting in distribution misalignment between the generated negatives and the model’s actual hallucination behavior, thereby offering limited improvement and failing to support adaptive and continual online learning [24]. To address these limitations, we propose a failure-guided negative generation strategy that directly targets self-generated hallucinated responses, enabling the real-time creation of more in-distribution counterexamples through text-to-image generation. Specifically, we sample and filter positive and negative response pairs from the model’s textual outputs. Then we employ LLMs to generate an image prompt based on the negative response, particular emphasizing its differences from the positive response, and subsequently synthesize the corresponding negative image using a diffusion model [25].
@@
Building upon this negative generation strategy, we further introduce an online vision-language preference learning framework (OViP) on both response and image sides, to dynamically construct and learn from preference data during training. Similar to reinforcement learning paradigms, OViP samples LVLMs’ outputs throughout the training process and creates both response-centric and image-centric preference pairs in real time. As illustrated in Figure 1, these dual signals supervise the model to generate outputs more faithfully grounded in visual content. By continuously sampling and integrating new preference pairs based on emerging failure patterns, OViP enables adaptive learning that aligns with the evolving output distributions of the LVLMs. This ongoing adaptation mitigates the limitations of static datasets and reduces the reliance on extensive manual curation, resulting in more robust and generalizable vision-language alignment.
We evaluate our framework across diverse multi-modal benchmarks, including hallucination-focused and general benchmarks. Through extensive experiments, we observe a notable trade-off between hallucination suppression and general language-vision capabilities, and some existing hallucination-elimination methods prove to overfit to hallucination-related metrics, achieving high scores without genuinely improving visual grounding. For example, some methods reduce hallucination by making descriptions or responses less informative. To address this, we refine existing evaluation protocols and introduce a more robust evaluation strategy that jointly assesses hallucination and visual-language competence. Experimental results demonstrate that our method consistently achieves improved performance on both the original and our refined metrics when applied to various LVLMs, demonstrating reduced hallucination and preserved general capability.
@@ 2 Methodology
Figure 2: Overview of OViP. Given an image and a query, we employ the current model $\pi_{t}$ to generate multiple responses, which are then evaluated by an external LLM with reference to the ground truth. We filter and select response pairs and use a diffusion model to generate corresponding negative images. The collected data are used to update $\pi_{t}$. The filtering strategy is detailed in Section 2.2.
@@ 2.1 Overview
Specifically, given an input image $\mathcal{I}^{+}$, an instruction $\mathcal{Q}$, and a reference response $\mathcal{A}^{*}$, OViP first samples multiple candidate responses using the target model $\pi$. These responses are then filtered and selected to form positive and negative pairs $(\mathcal{A}^{+},\mathcal{A}^{-})$. Based on the semantic discrepancies between the response pairs, contrastive images $\mathcal{I}^{-}$ are further synthesized to describing the negative responses. Finally, both image-level and response-level contrastive losses are applied to update the target model $\pi$. A detailed workflow of the OViP algorithm is provided in Appendix C.
@@ 2.2 In-Distribution Preference Data Construction
To effectively learn from preference data, it is crucial to construct pairs with sufficient contrast in quality [26], yet usually struggled when the generated responses exhibit similar quality. To address this, we dynamically construct preference pairs by selecting response pairs within each batch that display significant score disparities. Specifically, for each set of candidate responses $\{\mathcal{A}^{i}\}_{i=1}^{k}$ with corresponding rewards $\{r^{i}\}_{i=1}^{k}$, we compute the standard deviation $\sigma_{r}$ of the reward scores and select pairs $\left(\mathcal{A}^{+},\mathcal{A}^{-}\right)$ that satisfy:
$|r^{+}-r^{-}|>\max\left(\delta,2\sigma_{r}\right)$
where $\delta$ is a fixed lower-bound margin. This criterion ensures that only response pairs exhibiting substantial contrast in reward scores are selected, effectively emphasizing informative differences between success and failure responses. Additionally, we enforce quality constraints by requiring that the accepted positive responses meet a predefined quality criterion (i.e., $r^{+}>\tau_{\text{pos}}$), while rejected negative responses fall below a specified threshold (i.e., $r^{-}<\tau_{\text{neg}}$). In cases where all candidate responses collectively perform poorly, we leverage offline ground-truth answers $\mathcal{A}^{*}$ as positive responses to guide the model learning effectively, a practice reminiscent of the mixed-policy approach in [27].
@@
$|r^{+}-r^{-}|>\max(\delta,2\sigma_{r}),$
$r^{+}>\tau_{\mathrm{pos}},\ r^{-}<\tau_{\mathrm{neg}}\big{\}}$
@@ 2.3 Image- and Response-Side Preference Optimization
To effectively align both textual and visual modalities during training, we formulate a unified optimization framework that simultaneously considers response-level and image-level preference signals. The overall optimization objective consists of two complementary components. The first is the text DPO loss[14], which guides the model to learn response-level preferences conditioned on the input image and instruction:
$\mathcal{L}_{\mathrm{Text}}\left(\mathcal{A}^{+},\mathcal{A}^{-};\mathcal{I}^{+},\mathcal{Q}\right)=\log\sigma\left(\beta\cdot\left[\log\frac{\pi_{\theta}(\mathcal{A}^{+}|\mathcal{I}^{+},\mathcal{Q})}{\pi_{\mathrm{ref}}(\mathcal{A}^{+}|\mathcal{I}^{+},\mathcal{Q})}-\log\frac{\pi_{\theta}(\mathcal{A}^{-}|\mathcal{I}^{+},\mathcal{Q})}{\pi_{\mathrm{ref}}(\mathcal{A}^{-}|\mathcal{I}^{+},\mathcal{Q})}\right]\right)$ (4)
In addition to response-level alignment, we incorporate a contrastive objective focused on the visual input. By keeping the query and response fixed, the model is required to learn preferences solely from differences in the visual input. On top of this, to further ensure that the model’s output maintains a reasonable and smooth probability distribution, we introduce the image-free term $\pi_{\theta}(\mathcal{A}|\mathcal{Q})$ and implement the image-side loss as in [20]:
@@ 3.1 Experimental Setup
We conduct our experiments on LLaVA-1.5-7B-hf and LLaVA-1.5-13B-hf[5], with CLIP ViT-L-336px as the visual encoder and Vicuna-7b/13b as the backbone respectively. The training dataset, sourced from [28], consists of 8,730 samples and 4,013 distinct image–query combinations, including image description, question answering, and some yes/no questions. Both the 7B and 13B models are trained for a single epoch using a cosine learning rate schedule with a global batch size of 16. We use LoRA [29] with a rank of 256 and alpha of 512. We set $\beta=\beta_{1}=\beta_{2}=0.1$ in Eq. 4 and Eq. 5. Learning rates are 1e-6 for 7B model and 5e-7 for 13B model.
@@
We compare OViP with SFT, DPO [14], mDPO [19] and GRPO [30]. As the original versions of SFT, DPO and mDPO are offline methods, we additionally implement iterative DPO and GRPO to facilitate a more comprehensive comparison. Furthermore, we evaluate several prior works with publicly available model weights, including HA-DPO [31], HALVA [18], RLAIF-V [16] and OPA-DPO [28]. Among them, our OViP and OPA-DPO use the same original training data, which is a subset of the dataset used by RLAIF-V.
@@ 3.2 Evaluation Metrics
The evaluation benchmarks are categorized into two types: hallucination-related evaluation and out-of-domain general capability evaluation.
Hallucination-Related Evaluation. These benchmarks assess LVLMs’ hallucination as well as visual competence. We provide illustrative examples in Appendix A.2 to highlight the necessity for improving the metrics for MMHal, AMBER_gen and ObjectHal.
  • MMHal-Bench (MMHal) [32] is a model-evaluated question-answering benchmark covering 8 categories and 12 topics. While the original evaluation strategy uses GPT-4 to judge model responses, the use of a text-only model will introduce considerable judging-time hallucinations and errors (as pointed by [33]). We following [33] and use gpt-4o-2024-05-13 for evaluation.
  • AMBER generative (AMBgen[34] is a judging-model-free benchmark for the image description task, comprising 1,004 samples. The metric Chair measures the object-level hallucination rate as the average precision of objects mentioned in the model’s descriptions, while Cover indicates the recall of objects. We observe a noticeable trade-off between these two metrics across various methods, where improvements in one often come at the expense of the other. To provide a more balanced and overall assessment, we introduce a new F1 score calculated as the harmonic mean of Chair and Cover.
  • Object HalBench (ObjectHal) [35] evaluates object-level completeness and hallucination rates using off-the-shelf LLMs. The generation prompts are augmented from [15]. Chairr denotes the response-level hallucination rate. Like AMBER_gen, we also introduce an object-level F1 metric to comprehensively measures the balance between hallucination and object coverage. Objects extraction is performed using gpt-4o-2024-05-13.
  • Llava-Bench-in-the-Wild (LV) [4] evaluates models’ abilities in visual understanding, reasoning, and instruction following, using 60 open-ended questions grounded in 24 diverse images from real-world and abstract scenarios. The evaluation is conducted using gpt-4o-2024-05-13.
  • AMBER discriminative (AMBERdis[34] includes 14,216 “Yes/No” questions regarding objects in image. We use the F1-score as its metrics.
To aggregate performance across five hallucination-related benchmarks, we introduce the Hallucination Reduction Index (HRI) as a unified measure of overall improvement. Specifically, for each benchmark, we compute the performance change relative to the baseline and apply proportional rescaling to map the values approximately within the range $[0,2]$. The final HRI score is obtained by summing the rescaled improvements from all benchmarks. As a dynamic metric that depends on actual performance, detailed calculations for each specific HRI are provided in Appendix A.1.
Out-of-domain General Capability Evaluation. In addition to the trade-off between hallucination reduction and completeness, we also observe a pronounced trade-off between hallucination reduction and general visual capability, i.e., hallucination-targeted models often exhibit performance declines on non-hallucination benchmarks. To assess this impact, we further evaluate LVLMs’ capabilities on general benchmarks, including RealworldQA [36], TextVQA [37], CVBench [38], MMStar [39]. We aggregate the results across these benchmarks and compute the Accuracy Difference, serving as a unified metric to quantify overall performance variation after training.
@@ 3.3 Main Results
AMBgen MMHal ObjectHal LV AMBdis HRI General
@@
LLaVA-1.5-7B Baseline 7.1 50.0 65.01 1.90 51.38 72.40 57.20 85.5 - -
HA-DPO[31] 5.6 49.4 64.86 1.95 37.15 73.81 57.30 85.4 -1.52 -11.59-
HALVA[18] 5.7 52.9 67.78 2.12 43.40 76.01 58.60 86.5 -9.08 -7.36
RLAIF-V∗♯[16] 3.1 49.8 65.79 2.54 09.35 69.78 58.90 86.4 -1.37 -6.74
OPA-DPO∗♯[28] 2.4 45.2 61.79 2.78 06.37 63.26 64.80 86.7 -5.60 -11.82-
@@
LLaVA-1.5-13B Baseline 6.5 51.0 65.99 2.24 46.18 76.73 62.60 89.1 - -
HALVA[18] 6.0 52.2 67.12 2.45 35.07 77.75 61.70 90.0 -4.22 -5.45
OPA-DPO∗♯[28] 2.8 47.8 64.08 2.88 05.88 64.46 64.70 89.3 -7.05 -15.25-
@@
Table 1 presents results of OViP and other methods across multiple benchmarks on various LVLM backbones. OViP consistently achieves significant improvements across most primary metrics while effectively preserving model’s general visual capabilities, in contrast to most other methods that exhibit varying degrees of degradation in general benchmarks. Moreover, OViP’s performance gains become increasingly pronounced with larger LVLM backbones, highlighting its scalability.
We further observe that our OViP can continuously learn from itself to enhance performance, with capabilities improving progressively over additional training epochs, demonstrating the self-improving nature of its online learning approach. Notably, even with just one training epoch, OViP surpasses HALVA and the two-epoch GRPO, both of which utilize twice the training data, underscoring the exceptional training efficiency of our approach.
A critical phenomenon often overlooked in previous work [17, 28, 21, 40, 19, 15, 16] is that offline hallucination methods tend to make model outputs less informative, as evidenced by declined Cover metric in AMBgen. However, information omission is also a form of hallucination, especially in image captioning tasks, where omitting entities in the image is effectively equivalent to the model assuming that those entities do not exist. While many prior methods reduce conventional hallucinations, they inadvertently exacerbate omission-based hallucinations. In contrast, OViP achieves a more balanced trade-off between informativeness and hallucination reduction.

### v2

@@ Abstract
Large vision-language models (LVLMs) remain vulnerable to hallucination, often generating content misaligned with visual inputs. Although recent training-based approaches aim to mitigate hallucination, they typically rely on predefined or randomly edited negative samples that do not reflect actual model errors, thus limiting training efficacy. In this work, we propose an Online Vision-language Preference Learning (OViP) framework that dynamically constructs contrastive training data based on the model’s own hallucinated outputs. By identifying semantic differences between sampled response pairs and synthesizing negative images using a diffusion model, OViP generates more relevant supervision signals in real time. This failure-driven training enables adaptive alignment of both textual and visual preferences. Moreover, we refine existing evaluation protocols to better capture the trade-off between hallucination suppression and expressiveness. Experiments on hallucination and general benchmarks demonstrate that OViP not only reduces hallucinations while preserving core multi-modal capabilities, but also substantially improves training efficiency. Code is available at https://github.com/lsjlsj35/Online-Vision-Language-Preference-Learning-for-VLM-Hallucination.
footnotetext: Please contact us at shujuanliu24@m.fudan.edu.cn.
@@ 1 Introduction
Large vision-language models (LVLMs) (Alayrac et al., 2022; Chen et al., 2023; 2024a; Liu et al., 2023; 2024b) have demonstrated remarkable performance across a wide range of multi-modal tasks (Dai et al., 2023; Li et al., 2023a; Bai et al., 2023; Wang et al., 2024b) by integrating pre-trained visual encoders with large language models (LLMs) to process and generate language grounded in visual inputs. However, LVLMs continue to struggle with persistent hallucination issues (Li et al., 2023b; Bai et al., 2024), often exhibiting incorrect references to visual content (Liu et al., 2024a; Zhou et al., 2023; Bai et al., 2024). These errors manifest as misattributing object properties, describing nonexistent entities, or fabricating spatial relationships that do not align with the image. Such inconsistencies undermine the model’s faithfulness to the input and hinder further reasoning capabilities, significantly limiting the reliability of LVLMs in real-world applications.
@@
Recent success of Direct Preference Optimization (DPO) (Rafailov et al., 2023) in LLMs alignment has inspired the exploration of multi-modal DPO to mitigate hallucination in LVLMs (Yu et al., 2024a; b; Xie et al., 2024; Sarkar et al., 2024). However, early efforts directly extend the original DPO designs from LLMs to LVLMs by constructing preference pairs solely on textual responses given the same image input, primarily focusing on response-side preference optimization and showing limited effectiveness. Recent advancements incorporate additional preference pairs conditioned on varying image inputs while keeping the same response, optimizing both visual and textual preference optimization (Wang et al., 2024a; Wu et al., 2025; Fu et al., 2025). This paradigm provides a complementary training signal that encourages the model to attend more closely to visual content.
@@
However, prior work mainly relies on existing paired datasets (Wu et al., 2025) or expert-defined patterns to construct negative image inputs, using techniques such as random cropping (Wang et al., 2024a), noise disruption (Zhou et al., 2024a), object removal (Lu et al., 2025), or human/LLMs generated element-replaced response for image editing (Xie et al., 2024). These strategies are typically not explicitly tied to model failures, resulting in distribution misalignment between the generated negatives and the model’s actual hallucination behavior, thereby offering limited improvement and failing to support adaptive and continual online [^1] We adopt the LLM community’s convention of using “online” to denote “on-policy” in RL. learning (Guo et al., 2024). To address these limitations, we propose a failure-guided negative generation strategy that directly targets self-generated hallucinated responses, enabling the real-time creation of more in-distribution counterexamples through text-to-image generation. Specifically, we sample and filter positive and negative response pairs from the model’s textual outputs. Then we employ LLMs to generate an image prompt based on the negative response, particular emphasizing its differences from the positive response, and subsequently synthesize the corresponding negative image using a diffusion model (Zhang et al., 2023).
@@
Building upon this negative generation strategy, we further introduce an online vision-language preference learning framework (OViP) on both response and image sides, to dynamically construct and learn from preference data during training. Similar to reinforcement learning paradigms, OViP samples LVLMs’ outputs throughout the training process and creates both response-centric and image-centric preference pairs in real time. As illustrated in Figure 1, these dual signals supervise the model to generate outputs more faithfully grounded in visual content. By continuously sampling and integrating new preference pairs based on emerging failure patterns, OViP enables adaptive learning that aligns with the evolving output distributions of the LVLMs. This ongoing adaptation mitigates the limitations of static datasets and reduces the reliance on extensive manual curation
We evaluate our framework on diverse multi-modal benchmarks, covering both hallucination-focused and general tasks. Based on our experiments, we find a notable trade-off between hallucination suppression and general capability or informativeness (what we refer to as “implicit hallucination”). To address this, we refine existing evaluation protocols and reveal that many prior methods tend to overestimate their improvements. Experimental results show that OViP delivers significant advantages in both performance and efficiency. Furthermore, we investigate the role of online training and visual signals, as well as their interactions, in shaping training effectiveness.
@@ 2 Methodology
Figure 2: Overview of OViP. Given an image and a query, we employ the current model $\pi_{t}$ to generate multiple responses, which are then evaluated by an external LLM with reference to the ground truth. We filter and select response pairs and then generate corresponding negative images. The collected data are used to update $\pi_{t}$. The filtering strategy is detailed in Section $2.$
@@ 2.1 Overview
Specifically, given an input image $\mathcal{I}^{+}$, an instruction $\mathcal{Q}$, and a reference response $\mathcal{A}^{*}$, OViP first samples multiple candidate responses using the target model $\pi$. These responses are then filtered and selected to form positive and negative pairs $(\mathcal{A}^{+},\mathcal{A}^{-})$. Based on the semantic discrepancies between the response pairs, contrastive images $\mathcal{I}^{-}$ are further synthesized to describing the negative responses. Finally, both image-level and response-level contrastive losses are applied to update the target model $\pi$. A detailed workflow of the OViP algorithm is provided in Table 4.
@@ 2.2 In-Distribution Preference Data Construction
To effectively learn from preference data, it is crucial to construct pairs with sufficient contrast in quality (Yu et al., 2025). We dynamically construct preference pairs by selecting response pairs within each batch that display significant score disparities. Specifically, for each set of candidate responses $\{\mathcal{A}^{i}\}_{i=1}^{k}$ with corresponding rewards $\{r^{i}\}_{i=1}^{k}$, we compute the standard deviation $\sigma_{r}$ of the reward scores and select pairs $\left(\mathcal{A}^{+},\mathcal{A}^{-}\right)$ that satisfy $|r^{+}-r^{-}|>\max\left(\delta,2\sigma_{r}\right)$ where $\delta$ is a fixed lower-bound margin. This criterion ensures that only response pairs exhibiting substantial contrast in reward scores are selected, effectively emphasizing informative differences between success and failure responses. Additionally, we enforce quality constraints by requiring that the accepted positive responses meet a predefined quality criterion (i.e., $r^{+}>\tau_{\text{pos}}$), while rejected negative responses fall below a specified threshold (i.e., $r^{-}<\tau_{\text{neg}}$). In cases where all candidate responses collectively perform poorly, we leverage offline ground-truth answers $\mathcal{A}^{*}$ as positive responses to guide the model learning effectively, a practice reminiscent of the mixed-policy approach in Yan et al. (2025).
@@
$|r^{+}-r^{-}|>\max(\delta,2\sigma_{r}),r^{+}>\tau_{\mathrm{pos}},\ r^{-}<\tau_{\mathrm{neg}}\big\}$
@@ 2.3 Image- and Response-Side Preference Optimization
To effectively align both textual and visual modalities during training, we formulate a unified optimization framework that simultaneously considers response-level and image-level preference signals. The overall optimization objective consists of two complementary components. The first is the text DPO loss(Rafailov et al., 2023), which guides the model to learn response-level preferences conditioned on the input image and instruction:
$\mathcal{L}_{\mathrm{Text}}\left(\mathcal{A}^{+},\mathcal{A}^{-};\mathcal{I}^{+},\mathcal{Q}\right)=-\log\sigma\left(\beta\cdot\left[\log\frac{\pi_{\theta}(\mathcal{A}^{+}|\mathcal{I}^{+},\mathcal{Q})}{\pi_{\mathrm{ref}}(\mathcal{A}^{+}|\mathcal{I}^{+},\mathcal{Q})}-\log\frac{\pi_{\theta}(\mathcal{A}^{-}|\mathcal{I}^{+},\mathcal{Q})}{\pi_{\mathrm{ref}}(\mathcal{A}^{-}|\mathcal{I}^{+},\mathcal{Q})}\right]\right)$ (4)
In addition to response-level alignment, we incorporate a contrastive objective focused on the visual input. By keeping the query and response fixed, the model is required to learn preferences solely from differences in the visual input. On top of this, to further ensure that the model’s output maintains a reasonable and smooth probability distribution, we introduce the image-free term $\pi_{\theta}(\mathcal{A}|\mathcal{Q})$ and implement the image-side loss as in Wu et al. (2025):
@@ 3.1 Experimental Setup
We conduct our experiments on LLaVA-1.5-7B-hf and LLaVA-1.5-13B-hf (Liu et al., 2024b), with CLIP ViT-L-336px as the visual encoder and Vicuna-7b/13b as the backbone respectively. The training dataset, sourced from Yang et al. (2025), consists of 8,730 samples and 4,013 distinct image–query combinations, including image description, question answering, and some yes/no questions. We use LoRA (Hu et al., 2022) with a rank of 256 and alpha of 512. Other settings are listed in Appendix B.2
@@
We compare OViP with SFT, DPO (Rafailov et al., 2023), mDPO (Wang et al., 2024a) and GRPO (Shao et al., 2024). As the original versions of SFT, DPO and mDPO are offline methods, we additionally implement iterative DPO and GRPO to facilitate a more comprehensive comparison. Furthermore, we evaluate several prior works with publicly available model weights, including HA-DPO (Zhao et al., 2023), HALVA (Sarkar et al., 2024), RLAIF-V (Yu et al., 2024b) and OPA-DPO (Yang et al., 2025). Among them, our OViP and OPA-DPO use the same original training data, which is a subset of the dataset used by RLAIF-V.
@@ 3.2 Evaluation Metrics
We conduct evaluations on five hallucination-related and four general capability benchmarks to assess hallucination mitigation and overall capability degradation.
Hallucination-Related Evaluation.
We evaluate hallucination in LVLM outputs using MMHal-Bench (MMHal) (Sun et al., 2024), AMBER generative (AMBgen(Wang et al., 2023), Object HalBench (ObjectHal) (Rohrbach et al., 2018), Llava-Bench-in-the-Wild (LV) (Liu et al., 2023), and AMBER discriminative (AMBdis(Wang et al., 2023). Detailed descriptions of the datasets, evaluation procedures, and metrics are provided in Appendix A.2
Prior work has primarily focused on assessing the precision of model outputs, i.e., whether the generated content contains explicit hallucinations. However, this perspective often overlooks the completeness of the output: a model may omit relevant entities (especially in image description tasks), leading to what we term implicit hallucinations. We argue that both explicit and implicit hallucinations are critical for a faithful evaluation of model reliability. Building on this perspective and the observation of failure cases where existing benchmarks can be hacked, we refine the evaluation protocols and introduce an F1 score for AMBgen and ObjectHal to better capture the extent of hallucination in generated responses. Illustrative failure cases of prior evaluation strategies are presented in Appendix A.3.
To aggregate performance across five benchmarks, we introduce the Hallucination Reduction Index (HRI) as a unified measure of overall improvement. HRI is computed by summing the normalized improvements from each benchmark to obtain the overall relative gain. The detailed calculation of HRI and the discussion of its fairness are provided in Appendix A.1.
General Capability Evaluation
To assess the trade-off between hallucination mitigation and general visual capability, we evaluate the trained models on general benchmarks, including RealworldQA (xAI, 2024), TextVQA (Singh et al., 2019), CVBench (Tong et al., 2024), MMStar (Chen et al., 2024b). We aggregate the results across these benchmarks and compute the Accuracy Difference, serving as a unified metric to quantify overall performance variation after training.
@@ 3.3 Main Results
AMBgen MMHal ObjectHal LV AMBdis
[…]
General
@@
LLaVA-1.5-7B
Baseline 7.1 50.0 65.01 1.90 51.38 72.40 57.20 85.5 - -
HA-DPO 5.6 49.4 64.86 1.95 37.15 73.81 57.30 85.4 -1.52 -11.59-
HALVA 5.7 52.9 67.78 2.12 43.40 76.01 58.60 86.5 -9.08 -7.36
RLAIF-V∗♯ 3.1 49.8 65.79 2.54 09.35 69.78 58.90 86.4 -1.37 -6.74
OPA-DPO∗♯ 2.4 45.2 61.79 2.78 06.37 63.26 64.80 86.7 -5.60 -11.82-
@@
LLaVA-1.5-13B
Baseline 6.5 51.0 65.99 2.24 46.18 76.73 62.60 89.1 - -
HALVA 6.0 52.2 67.12 2.45 35.07 77.75 61.70 90.0 -4.22 -5.45
OPA-DPO∗♯ 2.8 47.8 64.08 2.88 05.88 64.46 64.70 89.3 -7.05 -15.25-
@@
Table 1 presents results for OViP and other methods across multiple benchmarks on various LVLM backbones. OViP consistently achieves significant improvements across most primary metrics while effectively preserving the model’s general visual capabilities (achieving +0.88 with one epoch for General AccDif and a slight drop of -1.01 for 2 epochs), whereas most other methods that exhibit varying degrees of degradation in general benchmarks. Moreover, OViP further improves with an additional training epoch. Notably, even with one epoch, OViP surpasses HALVA and 2-epoch GRPO, both of which utilize twice as much training data, but still yield lower HRI and suffer from general ability degradation.
A critical phenomenon often overlooked in previous work (Xie et al., 2024; Yang et al., 2025; Fu et al., 2025; Xiao et al., 2025; Wang et al., 2024a; Yu et al., 2024a; b) is that offline methods generally impair models’ general capability while also introducing implicit hallucinations (as discussed in subsection 3.2). This issue is particularly evident in OPADPO, where Chair score on AMBgen drops to 2.4, and Cover metric decreases from the initial 50.0 to 45.2, far below other methods. An illustrative example of such omission is in Figure 9 in Appendix. Moreover, excessive training further exacerbates this problem: as shown in Table 1, several DPO-like methods  (HA-DPO, HALVA, RLAIF-V, OPA-DPO) trained for more than two epochs suffer from much larger declines in general capability compared to DPO and mDPO trained for only one epoch. At the same time, except for HALVA, their HRI scores are also lower than those of DPO and mDPO, which mainly influenced by the low F1 scores on AMBgen and ObjectHal. With these possible signs of overfitting, we suggest that some improvements reported in prior work may be overestimated.